Graph the function f(x) = x + 2

Answer by jim_thompson5910(35256)

Plug in x = 0 to get f(0) = 0+2 = 2, so we have the point (0,2)

Plug in x = 1 to get f(1) = 1+2 = 3, so we have the point (1,3)

Now plot the two points and draw a straight line through the two points to graph f(x) = x+2